# GMG: A Video Prediction Method Based on Global Focus and Motion Guided

### v2

@@ Abstract
Recent years, weather forecasting has gained significant attention. However, accurately predicting weather remains a challenge due to the rapid variability of meteorological data and potential teleconnections. Current spatiotemporal forecasting models primarily rely on convolution operations or sliding windows for feature extraction. These methods are limited by the size of the convolutional kernel or sliding window, making it difficult to capture and identify potential teleconnection features in meteorological data. Additionally, weather data often involve non-rigid bodies, whose motion processes are accompanied by unpredictable deformations, further complicating the forecasting task. In this paper, we propose the GMG model to address these two core challenges. The Global Focus Module, a key component of our model, enhances the global receptive field, while the Motion Guided Module adapts to the growth or dissipation processes of non-rigid bodies. Through extensive evaluations, our method demonstrates competitive performance across various complex tasks, providing a novel approach to improving the predictive accuracy of complex spatiotemporal data.
Index Terms:
Spatio-temporal, Video Prediction, Motion Guided, Global receptive field.
@@ I Introduction
In recent years, spatiotemporal prediction has garnered unprecedented attention due to the growing demand for prediction tasks across various domains, particularly in weather forecasting [1, 2] and traffic prediction [3, 4]. Accurate predictions can bring significant social and economic value. Currently, spatiotemporal prediction models can be categorized into four main types: RNN, CNN, CNN-RNN, and ViT.
@@
CNN models focus on capturing multiscale features through techniques such as multiscale convolution stacking, upsampling, and skip connections. The strength of CNN models lies in their simplicity, allowing them to capture multiscale features efficiently, although they often struggle to capture long-term temporal features. Notable CNN models include PredCNN [10], SimVP [11], and TAU [12], which have demonstrated excellent performance and interpretability using streamlined predictive frameworks, offering diverse solutions for spatiotemporal sequence prediction.
@@
ViT primarily utilizes the Transformer architecture, which has gained popularity in recent years, leveraging the attention mechanism to capture image features. SwinLSTM [16], SimVP-ViT [17], and PredFormer [18] are built on Vision Transformer [19] or Swin Transformer [20] backbones. As an emerging approach, ViT models have been proven to possess strong predictive capabilities. However, due to the high computational complexity of attention mechanisms, they typically require more training resources.
Meteorological data prediction holds significant research value due to its timeliness and immense societal impact. However, spatiotemporal data often exhibit complex non-stationary processes with rapid short-term variations, making prediction tasks highly challenging. This paper focuses on two core issues (Figure 1): deformation caused by object motion and potential long-range correlations.
@@
Figure 1: The two main real-world challenges addressed in this paper. Position and shape changes of rainfall regions over time: For example, in region B, while moving eastward, the north-south extent of the rainfall gradually expands, adding complexity to prediction tasks. Long-range correlations: In real-life scenarios, a traffic jam in region C may impact traffic in another part of the city, such as region D. These long-range correlations are often difficult to capture, limiting the accuracy of predictions for such data.
@@
Taking precipitation data as an example, rain clouds not only shift in a specific direction but also undergo overall deformation. Moreover, the dissipation and growth of clouds over time are particularly difficult to predict accurately. Similarly, in real life, a traffic jam on one street may impact another distant street. In meteorological data, analogous phenomena exist where climate anomalies exhibit correlations over long distances. However, the complexity and unpredictability of such long-range correlations significantly limit prediction accuracy.
@@
Previous studies have developed models that decouple motion patterns to better predict object motion. MotionRNN, for instance, decomposes motion into transient variation and trending momentum to model object motion dynamics. MAU [21] separately trains augmented motion information (AMI) and current spatial state to capture reliable inter-frame motion information. MMVP [22] constructs an appearance-independent motion matrix to decouple motion and appearance information, inferring future object motion from image frames while maintaining appearance consistency across frames. These studies provide valuable theoretical foundations. Building on these, we propose a Motion Guided Module (MGM) to better predict deformations arising during object motion. Extending MotionRNN’s displacement decomposition framework, we introduce a deformation pattern regulated by a balance factor $\alpha$ and a decay factor $\beta$ to simulate this non-stationary process effectively.
@@
Despite the advances achieved by current methods through enhanced information flow mechanisms, multiscale convolution stacking, or sliding-window strategies to broaden the model’s global perspective, a critical issue remains: whether based on convolutions or sliding windows, the receptive field of these approaches is inherently limited. This limitation makes it challenging for models to fully capture and generalize global features when processing spatiotemporal data with potential long-range correlations. To address this, we propose a Global Focus Module (GFM) that enhances the model’s global perspective by extracting global features from the input image $X$ and integrating them with the hidden state $H$.
@@
The proposed GMG predictive model consists of four main components: SpatioTemporal ConvLSTM (ST-ConvLSTM), Motion Guided Module (MGM), Self-Attention Memory, and Global Focus Module (GFM). We aim to offer a general solution for video prediction, particularly for meteorological data, and provide a novel approach to achieving accurate meteorological forecasting.The main contributions of this paper are summarized as follows:
@@
  • We propose a more comprehensive object motion decoupling method, the Motion Guided Module (MGM), which captures non-elastic deformation characteristics by introducing a balance factor and a decay factor.
@@
  • By integrating MGM and GFM, we propose the GMG video prediction framework. Extensive experiments on five different types of datasets demonstrate that GMG achieves state-of-the-art performance, providing an innovative solution for video prediction.
@@ II Related Work
Over the past decade, spatiotemporal prediction using recurrent structures has originated from ConvLSTM, a simple yet effective framework that remains widely used today. Following this, numerous RNN-based models have been proposed. PredRNN [4] introduced ST-ConvLSTM, which added a Spatiotemporal Memory module parallel to the cell state in ConvLSTM and organized a cross-layer information flow. This design allowed the model’s final output to comprehensively capture hierarchical spatiotemporal features. Building upon PredRNN, PredRNN++ [5] regulated the sequence of spatiotemporal memory and cell states, leading to the introduction of Causal LSTM. Additionally, the authors proposed the Gradient Highway Unit, enabling rapid transmission of hidden states between adjacent basic units to mitigate gradient vanishing along the temporal dimension.
@@ III Methods
III-A Overall Architecture
The overall architecture of GMG is presented on the left side of Fig. 2. The input image $X$ is processed through ST-ConvLSTM to update the hidden state $H_{t}$. Subsequently, $H_{t}$ is fed into the Global Focus Module (GFM), where it is fused with global features to produce $H_{t}^{g}$, enabling the model to perceive global characteristics of the input data. The primary role of SAM [24] is to integrate $H_{t}^{g}$ with the spatiotemporal memory $M_{t}$ to capture spatial contextual information, thereby enhancing the model’s ability to maintain long-term feature dependencies. Finally, $\hat{H}_{t}^{g}$ is input into the Motion Guided Module (MGM), which regulates the model’s motion state learning ability through two state factors, $\alpha$ and $\beta$. GMG retains the cross-layer transmission of spatiotemporal memory and the Gradient Highway Unit [5], ensuring effective utilization of various types of valuable information.
III-B Global Focus Module
@@
First, we apply Adaptive Average Pooling (AAP) to extract the Global Features ($\bm{G_{F}}\in\mathbb{R}^{T\times C\times H\times W}$) from the input image:
@@
$\bm{X_{t}^{g}}=\texttt{ReLU}(\texttt{Linear}(\bm{G_{F}})).$ (2)
@@
$\mathcal{\bm{G}}=\texttt{Sigmoid}(\texttt{Conv2d}(\texttt{Concat}[\bm{H_{t}},\bm{X_{t}^{g}}])),$ (3)
@@
Where, $\bm{X_{t}^{g}},\bm{H_{t}}\in\mathbb{R}^{T\times C\times H\times W}$, $\circ$ denote the Hadamard product. To ensure that the updated hidden state includes more global information across different scales, we apply a multi-scale focus method:
@@
III-C Motion Guided Module
Inspired by the motion decomposition in MotionGRU [7], we propose a decomposition model based on motion evolution. This model divides the spatial movement of objects into transient variations $\mathcal{F}$ and trending momentum $\mathcal{D}$. $\mathcal{F},\mathcal{D}\in\mathbb{R}^{T\times 2k^{2}\times\frac{H}{2}\times\frac{W}{2}}$, $k$ is filter size. Additionally, it generalizes non-displacement changes of objects as the cumulative effects of two physical processes: collision and dissipation. To effectively capture the features of object deformation, we introduce a balance factor $\bm{\alpha}$ and a decay factor $\bm{\beta}$, which illustrated on the right side of Fig. 2.

### v3

@@ Abstract
Recent years, video prediction has gained significant attention particularly in weather forecasting. However, accurately predicting weather remains a challenge due to the rapid variability of meteorological data and potential teleconnections. Current spatiotemporal forecasting models primarily rely on convolution operations or sliding windows for feature extraction. These methods are limited by the size of the convolutional kernel or sliding window, making it difficult to capture and identify potential teleconnection features in meteorological data. Additionally, weather data often involve non-rigid bodies, whose motion processes are accompanied by unpredictable deformations, further complicating the forecasting task. In this paper, we propose the GMG model to address these two core challenges. The Global Focus Module, a key component of our model, enhances the global receptive field, while the Motion Guided Module adapts to the growth or dissipation processes of non-rigid bodies. Through extensive evaluations, our method demonstrates competitive performance across various complex tasks, providing a novel approach to improving the predictive accuracy of complex spatiotemporal data.
@@ I Introduction
In recent years, spatiotemporal prediction has garnered unprecedented attention due to the growing demand for prediction tasks across various domains, particularly in weather forecasting [1, 2, 3] and traffic prediction [4, 5]. Accurate predictions can bring significant social and economic value. Currently, spatiotemporal prediction models based on deep learning method can be categorized into four main types: RNN, CNN, CNN-RNN, and ViT.
@@
CNN models focus on capturing multiscale features through techniques such as multiscale convolution stacking, upsampling, and skip connections. The strength of CNN models lies in their simplicity, allowing them to capture multiscale features efficiently, although they often struggle to capture long-term temporal features. Notable CNN models include PredCNN [11], SimVP [12], TAU [13] and TrajectoryCNN [15], which have demonstrated excellent performance and interpretability using streamlined predictive frameworks, offering diverse solutions for spatiotemporal sequence prediction.
@@
ViT primarily utilizes the Transformer architecture, which has gained popularity in recent years, leveraging the attention mechanism to capture image features. SwinLSTM [18], SimVP-ViT [19] and PredFormer [20] are built on Vision Transformer [21] or Swin Transformer [22] backbones. As an emerging approach, ViT models have been proven to possess strong predictive capabilities. However, due to the high computational complexity of attention mechanisms, they typically require more training resources.
Given its timeliness and societal relevance, meteorological data prediction is of great research interest. Yet, spatiotemporal models, particularly convolution-based ones for feature learning, encounter two challenges (Fig. 1): (1) Localized feature extraction limits the ability to capture long-range dependencies, crucial for modeling large-scale meteorological patterns and non-local effects such as teleconnections. (2) The non-rigid nature of atmospheric elements (e.g., cloud formations and storm systems) makes it difficult to accurately model their evolution over time.
@@
Figure 1: The two main real-world challenges addressed in this paper. Position and shape changes of rainfall regions over time: For example, in region B, while moving eastward, the north-south extent of the rainfall gradually expands, adding complexity to prediction tasks. Long-range correlations: In real-life scenarios, a traffic jam in region C may be related to traffic in another part of the city, such as region D. These long-range correlations are often difficult to capture, limiting the accuracy of predictions for such data.
@@
Taking precipitation data as an example, rain clouds not only shift in a specific direction but also undergo overall deformation. The dissipation and growth of clouds over time are particularly difficult to predict accurately. In addition, in real life, a traffic jam on one street may impact another distant street. In meteorological data, analogous phenomena exist where climate anomalies exhibit correlations over long distances. However, the complexity and unpredictability of such long-range correlations significantly limit prediction accuracy.
@@
Previous studies have developed models that decouple motion patterns to better predict object motion. MotionRNN, for instance, decomposes motion into transient variation and trending momentum to model object motion dynamics. MAU [23] separately trains augmented motion information (AMI) and current spatial state to capture reliable inter-frame motion information. MMVP [24] constructs an appearance-independent motion matrix to decouple motion and appearance information, inferring future object motion from image frames while maintaining appearance consistency across frames. MAFE [26] proposes a Motion Perceptual Loss and extracts object-perceptive features through an attention module. HO-LMC [27] proposes a high-order motion encoder to capture high-order motion dynamics in image sequences. These studies provide valuable theoretical foundations. However, there is still a lack of effective feature extraction methods for non-rigid object motion.
Inspired by non-rigid motion modeling, this paper introduces a Motion Guided Module (MGM). Building upon conventional motion-state decomposition, MGM incorporates two deformation factors—balance factor $\alpha$ and decay factor $\beta$—to characterize local growth and dissipation, global deformation, and inelastic morphological changes of a target during motion. The balance factor regulates the contribution of local motion to the overall shape, ensuring a spatially coherent distribution of motion information. The attenuation factor models the temporal decay or dissipation of morphology, thereby enabling more precise prediction of dynamic changes in non-rigid targets. Through this design, MGM captures fine-grained features of complex target motions and improves the prediction accuracy of non-rigid motion sequences.
@@
Meanwhile, existing approaches extend a model’s global receptive field through enhanced information-flow mechanisms, multi-scale convolutional stacks, or sliding-window strategies. However, a key challenge remains: whether based on convolution or sliding windows, the receptive field is inherently limited, making it difficult for models to fully capture and generalize global features when dealing with spatiotemporal data with potential long-range dependencies. To address this issue, this paper proposes a Global Focus Module (GFM), a non-attention, low-complexity global feature aggregation method. By extracting global features from the input image $X$ and integrating them with the hidden state $H$, GFM substantially enhances the model’s global perspective. Compared with traditional attention-based global modeling methods, GFM can capture latent long-range dependencies without significantly increasing computational complexity, thereby improving the model’s predictive capability on large-scale spatiotemporal sequences.
@@
The proposed GMG (Global focus and Motion Guided) predictive model consists of four main components: SpatioTemporal ConvLSTM (ST-ConvLSTM), Motion Guided Module (MGM), Self-Attention Memory, and Global Focus Module (GFM). We aim to offer a general solution for video prediction, particularly for meteorological data, and provide a novel approach to achieving accurate meteorological forecasting. The main contributions of this paper are summarized as follows:
@@
  • We propose a more comprehensive object motion decoupling method, the Motion Guided Module (MGM), which captures non-rigid deformation characteristics by introducing a balance factor and a decay factor.
@@
  • By integrating MGM and GFM, we propose the GMG video prediction framework. Extensive experiments on six different types of datasets demonstrate that GMG achieves state-of-the-art performance, providing an innovative solution for video prediction.
@@ II Related Work
Over the past decade, spatiotemporal prediction using recurrent structures has originated from ConvLSTM, a simple yet effective framework that remains widely used today. Following this, numerous RNN-based models have been proposed. PredRNN [5] introduced ST-ConvLSTM, which added a Spatiotemporal Memory module parallel to the cell state in ConvLSTM and organized a cross-layer information flow. This design allowed the model’s final output to comprehensively capture hierarchical spatiotemporal features. Building upon PredRNN, PredRNN++ [6] regulated the sequence of spatiotemporal memory and cell states, leading to the formation of Causal LSTM. Additionally, PredRNN++ incorporates the Gradient Highway Unit (GHU), which facilitates fast information flow between adjacent recurrent units and effectively mitigates the vanishing gradient problem along the temporal dimension.
@@ III Methods
In this section, we introduce the proposed GMG prediction model. As mentioned above, our model is designed to address two key challenges in spatiotemporal forecasting:
(1) How to effectively capture non-rigid motion deformations to improve the accuracy of object movement prediction?
(2) How to enhance the model’s ability to extract global dependencies in order to better capture long-range correlations (eg., teleconnections) in spatiotemporal data?
In the following subsections, we first define the problem and then introduce in detail how our GMG model addresses the two aforementioned challenges.
III-A Problem Definition
The video prediction problem can be formulated as follows: Given a sequence of consecutive images $X$ over a certain period, the goal is to find a set of parameters $\theta^{*}$ that maximize the probability of the predicted image sequence $Y$:
$\theta^{*}=\arg\max_{\theta}P(Y|X;\theta).$ (1)
where $\theta$ represents the model parameters to be trained.
In this work, we adopt the $L_{2}$ loss for optimization, which assumes the prediction error follows a Gaussian distribution. Under this assumption, maximizing the likelihood is equivalent to minimizing the following loss:
$\mathcal{L}_{L_{2}}=\sum_{i}(Y_{i}-\hat{Y}_{i})^{2},$ (2)
where $\hat{Y}=F_{\theta}(X)$ represents the predicted future frames generated by the model $F_{\theta}$ given the input sequence $X$. The $L_{2}$ loss penalizes larger deviations more strongly, encouraging smooth and consistent predictions.
III-B Overall Architecture
The overall architecture of GMG is presented on the left side of Fig. 2. The input image $X$ is processed through ST-ConvLSTM to update the hidden state $H_{t}$. Subsequently, $H_{t}$ is fed into the Global Focus Module (GFM), where it is fused with global features to produce $H_{t}^{g}$, enabling the model to perceive global characteristics of the input data. The primary role of Self Attention Memory (SAM) from SA-ConvLSTM [28] is to integrate $H_{t}^{g}$ with the spatiotemporal memory $M_{t}$ to capture spatial contextual information, thereby enhancing the model’s ability to maintain long-term feature dependencies. Finally, $\hat{H}_{t}^{g}$ is input into the Motion Guided Module (MGM), which regulates the model’s motion state learning ability through two state factors, $\alpha$ and $\beta$. GMG retains the cross-layer transmission of spatiotemporal memory and the Gradient Highway Unit [6], ensuring effective utilization of various types of valuable information.
III-C Global Focus Module
@@
First, we apply Adaptive Average Pooling (AAP) to extract the Global Features ($\bm{G_{F}}$) from the input image $\bm{X_{t}}\in\mathbb{R}^{T\times C\times H\times W}$:
@@
Adaptive Average Pooling (AAP) divides the input feature map $X_{t}$ into rectangular regions and outputs the average value of all elements within each region. Formally, for the $i$-th feature map and region $R_{j}$:
$\mathcal{AAP}(\bm{X_{t}})_{i,j}=\frac{1}{|R_{j}|}\sum_{k\in R_{j}}\bm{X_{t}}^{(i)}(k)$ (4)
where $\bm{X_{t}}^{(i)}(k)$ is the element at position $k$ in region $R_{j}$ of the $i$-th feature map, $AAP(X_{t})_{i,j}$ denotes the result of average pooling over the $j$-th region of the $i$-th channel feature map at time $t$. $|R_{j}|$ is the number of elements in $R_{j}$. In the actual implementation, the output size is set to $1\times 1$, which $\bm{G_{F}}\in\mathbb{R}^{T\times C\times 1\times 1}$.
@@
$\bm{X_{t}^{g}}=\texttt{ReLU}(\texttt{Linear}(Enc(\bm{G_{F}}))).$ (5)
where, $Enc(\cdot)$ maps $\bm{G_{F}}$ from the shape $\mathbb{R}^{T\times C\times 1\times 1}$ to $\mathbb{R}^{T\times C\times H\times W}$.
@@
$\mathcal{\bm{G}}=\texttt{Sigmoid}(\texttt{Conv2d}_{W_{\mathcal{\bm{G}}}}(\texttt{Concat}[\bm{H_{t}},\bm{X_{t}^{g}}])),$ (6)
@@
Where, $\bm{X_{t}^{g}},\bm{H_{t}}\in\mathbb{R}^{T\times C\times H\times W}$, $W_{\mathcal{\bm{G}}}$ represents convolution kernel, $\circ$ denote the Hadamard product. To ensure that the updated hidden state includes more global information across different scales, we apply a multi-scale focus method:
@@
III-D Self-Attention Memory
To further enhance the hidden state $\bm{H_{t}^{g}}$ after global feature correction, we introduce the Self-Attention Memory (SAM) module, which captures the relationship between the current hidden state $\bm{H_{t}^{g}}$ and the long-term memory $\bm{M_{t-1}}$ through a self-attention mechanism.
The SAM takes two inputs: the globally enhanced hidden state $\bm{H_{t}^{g}}\in\mathbb{R}^{T\times C\times H\times W}$ and the memory from the previous timestep $\bm{M_{t-1}}\in\mathbb{R}^{T\times C\times H\times W}$. To facilitate attention computation, we first flatten the spatial dimensions into a sequence. Let $N=H\times W$, and reshape the inputs as $\bm{H_{t}^{g}},\bm{M_{t-1}}\in\mathbb{R}^{T\times C\times N}$.
Next, the hidden state $\bm{H_{t}^{g}}$ is linearly projected to obtain the query $\bm{Q_{h}}\in\mathbb{R}^{T\times d\times N}$, and the memory $\bm{M_{t-1}}$ is projected to generate the key and value, $\bm{K_{m}},\bm{V_{m}}\in\mathbb{R}^{T\times d\times N}$, where $d$ is the attention embedding dimension. The similarity score between query and key is computed as:
$\bm{e_{m}}=\bm{Q_{h}}^{\top}\bm{K_{m}}\in\mathbb{R}^{T\times N\times N},$ (10)
This is followed by a softmax operation to obtain normalized attention weights:
$s_{m;i,j}=\frac{\exp(e_{m;i,j})}{\sum_{k=1}^{N}\exp(e_{m;i,k})},$ (11)
The memory features are then aggregated using these weights:
$\bm{Z_{m;i}}=\sum_{j=1}^{N}s_{m;i,j}\cdot\bm{V_{m;j}}=\sum_{j=1}^{N}s_{m;i,j}\cdot\bm{W_{mv}}\bm{M_{t-1;j}},$ (12)
where $\bm{W_{mv}}$ is a learnable projection matrix, and $\bm{M_{t-1;j}}$ denotes the $j$-th spatial location of the memory.
Finally, the memory-based output $\bm{Z_{m}}$ is concatenated with a self-attention output $\bm{Z_{h}}$ computed from $\bm{H_{t}^{g}}$, and the combined result is fused via a linear projection:
$\bm{Z}=\bm{W_{z}}[\bm{Z_{h}};\bm{Z_{m}}],$ (13)
where $\bm{W_{z}}$ is a learnable fusion matrix. This fused representation $\bm{Z}$ is then passed to a gated memory update mechanism described in the following section.
To maintain long-term consistency, the memory is updated through a gated mechanism inspired by LSTM, but optimized for efficiency. Instead of full convolution, we use depthwise separable convolutions [37] to reduce parameters:
$\bm{i_{t}^{a}}$ $=\sigma({W_{zi}}*\bm{Z}+{W_{hi}}*\bm{H_{t}^{g}}+\bm{b_{i}}^{a}),$ (14)
$\bm{C_{t}^{a}}$ $=\tanh({W_{zc}}*\bm{Z}+{W_{hc}}*\bm{H_{t}^{g}}+\bm{b_{c}}^{a}),$ (15)
$\bm{M_{t}}$ $=(1-\bm{i_{t}^{a}})\circ\bm{M_{t-1}}+\bm{i_{t}^{a}}\circ\bm{C_{t}^{a}}.$ (16)
The final output hidden state $\hat{\bm{H_{t}}}$ is modulated by an output gate:
$\bm{o_{t}^{a}}$ $=\sigma({W_{zo}}*\bm{Z}+{W_{ho}}*\bm{H_{t}^{g}}+\bm{b_{o}^{a}}),$ (17)
$\bm{\hat{{H_{t}^{g}}}}$ $=\bm{o_{t}^{a}}\circ\bm{M_{t}}.$ (18)
Overall, the SAM module enables efficient retrieval and integration of historical contextual information through attention-weighted memory and lightweight gating mechanisms. Its use of depthwise separable convolutions makes it suitable for deployment in high-resolution scenarios.
It is worth noting that SAM and GFM are complementary. Unlike conventional global attention, GFM does not compute pairwise attention weights across regions; instead, it first extracts global statistical features via adaptive pooling and then applies gated multi-scale convolutions to globally calibrate the hidden state, thereby explicitly introducing long-range dependencies and cross-region correlations while avoiding high computational cost and memory overhead. Building on this, SAM retrieves and integrates long-term memory on demand along the temporal dimension, performing its self-attention operation only on compressed memory representations and employing depthwise separable convolutions for efficient gated updates. Consequently, the combination of GFM and SAM rapidly focuses on global spatial context while efficiently leveraging historical information over time, enabling richer and more stable spatiotemporal representations at lower complexity than global attention mechanisms—particularly well suited for video prediction tasks involving complex dynamics and long time horizons.
III-E Motion Guided Module
Inspired by the motion decomposition in MotionGRU [8], we propose a decomposition model based on motion evolution. This model divides the spatial movement of objects into transient variations $\mathcal{F}$ and trending momentum $\mathcal{D}$. $\mathcal{F},\mathcal{D}\in\mathbb{R}^{T\times 2k^{2}\times\frac{H}{2}\times\frac{W}{2}}$, $k$ is filter size. Additionally, MGM introduces for the first time in video prediction two physically interpretable deformation factors—the balance factor and the decay factor (Fig. 2). The balance factor models the “growth trend” of motion, reflecting a target’s tendency to preserve momentum and internal deformation during movement, thereby allowing the model to account for both instantaneous changes and historical motion states during prediction. The decay factor simulates the temporal dissipation of motion energy, applying an exponential decay to motion trends to suppress the influence of outdated dynamics, thus enhancing stability when handling long video sequences.